# TEXasGAN: Tactile Texture Exploration and Synthesis System Using Generative Adversarial Network

### v1

@@ VI Discussion
Therefore, in future work, we plan to explore other generative models to enhance the diversity and precision of the generated vibrations. For example, incorporating diffusion models could help improve the variety and fidelity of the output [35]. Diffusion models can be super easily to train compared with GAN and can provide high quality results. Additionally, the current optimization process is quite time consuming and requires subjects to focus intensely on detecting subtle differences while adjusting the slider position. Streamlining the interface for more intuitive operation could significantly improve user experience. To further create realistic interaction experiences, we will consider using devices that better simulate texture perception, integrating both spatial and temporal stimulation.

### v2

@@ VI Discussion
Therefore, in future work, we plan to explore other generative models to enhance the diversity and precision of the generated vibrations. For example, incorporating diffusion models could help improve the variety and fidelity of the output [35]. Diffusion models can be easily to train compared with GAN and can provide high quality results. Additionally, the current optimization process is quite time consuming and requires subjects to focus intensely on detecting subtle differences while adjusting the slider position. Streamlining the interface for more intuitive operation could significantly improve user experience. To further create realistic interaction experiences, we will consider using devices that better simulate texture perception, integrating both spatial and temporal stimulation.